OpenStudy (anonymous):

10 , 19 , ?, 32 , 40 Find ?

OpenStudy (anonymous):

@ajprincess @RaphaelFilgueiras

OpenStudy (anonymous):

i think it is 27.5.

Parth (parthkohli):

@Yahoo! can you explain? :/

OpenStudy (anonymous):

i dont know the real answer

Parth (parthkohli):

The only thing I can find is that there's a difference of 9 in the first two and then 8 in the last two :/

OpenStudy (anonymous):

first 9 is added next it shoulb be 8.5 then 8

Parth (parthkohli):

That's what I am thinking but I guess that the thing should be an integer

OpenStudy (anonymous):

yup

Parth (parthkohli):

Any idea, cwrw?

Parth (parthkohli):

Expee?

Parth (parthkohli):

lol

Parth (parthkohli):

How do you derive the sequence?

OpenStudy (anonymous):

10, 19, 29, 32, 40 reason>>>

OpenStudy (anonymous):

10, 19, 22, 32, 40 reson/////////////confuseedddd

Parth (parthkohli):

It says 22 in the other one. There are 2 results :/

OpenStudy (anonymous):

Confused??

OpenStudy (unklerhaukus):

can i ask where this question is coming from?

Parth (parthkohli):

Yahoo! Answers lol :P

OpenStudy (unklerhaukus):

i dont think there is a good answer to 10 , 19 , ?, 32 , 40 Find ?

OpenStudy (unklerhaukus):

unless is it 25.5

OpenStudy (ajprincess):

I think it's 26. Though nt sure. 10+9=19 19+7=26 26+6=32 32+8=40 9-7=2 7-6=1 8-6=2

OpenStudy (maheshmeghwal9):

I think no correct answer until the sequence has some more values:/

OpenStudy (maheshmeghwal9):

becoz all r getting different answers but from their point of view all r correct.

Parth (parthkohli):

If we'd have had one value added, we'd have recognized the patter :)

OpenStudy (maheshmeghwal9):

ya pattern*

OpenStudy (anonymous):

20 minutes later...mind=blown